# Promise and Peril: Stellar Contamination and Strict Limits on the Atmosphere Composition of TRAPPIST-1c from JWST NIRISS Transmission Spectra

### v1

@@ 5.4 Implications for the Presence of an Atmosphere on TRAPPIST-1c
Ultimately, our results from the analysis of the observed transmission spectrum of TRAPPIST-1 c are in line with modeling predictions that it would not have been able to retain any substantial atmosphere over the 5–9 Gyr lifetime of its host star (Burgasser & Mamajek, 2017), regardless of its initial volatile content (Krissansen-Totton, 2023).
@@
The thin, 1 bar atmospheres that remain allowed by our observations, even if they correspond to high mean molecular weights, are unlikely to be retained on TRAPPIST-1 c from a theoretical perspective. Even if Jeans escape (generally less efficient than blow-off) is expected to dominate for high-mean molecular weight atmospheres atop TRAPPIST-1 c, the escape rates even for resilient N2-CO2 atmospheres range from one Earth atmosphere per Myr to one Earth atmosphere mass per Gyr (Van Looveren et al., 2024) at the present-day irradiation of this planet. The implications are even stricter when considering the fact that any atmosphere on TRAPPIST-1 c also would have had to withstand the early XUV-active stages of the late M dwarf’s evolution (Bolmont et al., 2017), and its position inside of the runaway greenhouse limit (Turbet et al., 2020) indicates that water could not have been efficiently shielded from escape via condensation in a liquid water ocean, even if this limit is pushed to closer host star-planet distances around M dwarfs in 3D models (Turbet et al., 2023).
@@ 6 Conclusions
Here, we presented the first JWST transmission spectrum of the second innermost planet in the TRAPPIST-1 system; TRAPPIST-1 c. Our 0.6 – 2.85 µm NIRISS/SOSS spectrum reveals strong contamination by the TLSE in both visits. However, unlike in previous multi-visit analyses of rocky planets around M dwarfs (e.g., May et al., 2023; Moran et al., 2023; Cadieux et al., 2024), including TRAPPIST-1 b (Lim et al., 2023), we find consistent stellar heterogeneity properties between our two visits, resulting in nearly identical manifestations of the TLSE. Our results support the need for joint modelling of the TLSE and atmosphere properties to achieve robust constraints on rocky, M dwarf planets in the era of JWST. Importantly, further stellar model fidelity is critically required to accurately model out the impact of the TLSE and reach the sensitivity to detect or rule out thin atmospheres on small planets via their transit spectra.
When accounting for stellar contamination in our atmosphere analysis, we find that thick, clear hydrogen-dominated compositions are ruled out at better than 3- significance, and even high-mean molecular weight compositions rich in H2O, NH3 or CO are excluded for an atmosphere thickness of 1 bar or more at better than 2. The NIRISS/SOSS transmission spectrum of TRAPPIST-1 c alone cannot rule out all high mean molecular weight compositions, especially for CH4- and CO2-rich cases, to which our spectrum is not highly sensitive. However, in light of modelling predictions that CH4-rich compositions are unlikely as CH4 is sensitive to photochemical destruction (Turbet et al., 2018), and considering that Jeans escape would quickly strip TRAPPIST-1 c of even CO2-rich atmospheres (Van Looveren et al., 2024), we conclude that this planet is unlikely to bear any atmosphere whatsoever.
@@ Appendix A NAMELESS Reduction and ExoTEP Light Curve Fitting
We reduce our two JWST NIRISS/SOSS visits of TRAPPIST-1c, starting from the uncalibrated raw data up to the extraction of the stellar spectra, using the NAMELESS pipeline (Coulombe et al., 2023; Feinstein et al., 2023). We first go through the Stages 1 and 2 steps of the STScI jwst pipeline (v1.12.5) (Bushouse et al., 2022), which include super-bias subtraction, reference pixel correction, non-linearity correction, ramp-fitting, and flat-fielding. We then proceed with a sequence of custom steps to correct for remaining sources of noise such as bad pixels, non-uniform background, cosmic-rays, and 1/ following the methods described in Benneke et al. (2024). We flag bad pixels by looking for pixels whose second derivative are significant outliers in the spatial domain. The counts of all flagged pixels are then corrected using the bicubic interpolation method of the scipy.interpolate.griddata function. We correct for the non-uniform background by individually scaling the two regions of the STScI model background333https://jwst-docs.stsci.edu/ that are separated by the sudden jump in background flux situated around spectral pixel x700. Any remaining cosmic rays are corrected by clipping any count that is more than four standard deviations away from the running median for all pixels. As for the 1/ noise correction, we follow the same method described in Coulombe et al. (2023) and Benneke et al. (2024) where we scale individually each column of the first and second spectral orders considering only pixels that are within a 30-pixel distance from the center of the traces to compute the 1/ noise. We then extract the stellar spectra from both spectral orders using a simple box aperture with a width of 36 pixels.
@@
We perform the broadband and spectroscopic fits of the light curves produced by the NAMELESS reduction with the ExoTEP framework (Benneke et al., 2019a, b). For our fit of the white-light curve, which is obtained by summing all wavelengths of the first spectral order (0.85–2.85 m), we keep free the mid-transit time (, considering uniform priors spanning the time of the observations for each visit), planet-to-star radius ratio (, [0.01,0.2]), semi-major axis (, [15,45]), and the impact parameter (, [0,1]). We consider the quadratic limb-darkening law and fit directly for and assuming large uniform priors ([-3,3]) following Coulombe et al. (2024). For the first visit, we consider a quadratic trend for the systematics model. We also include in the fit a Matérn 3/2 Gaussian Process using the celerite python (Foreman-Mackey et al., 2017) to remove higher-frequency trends that are not corrected by our systematics model. The timescale and amplitude of the GP model are kept free assuming large uninformative uniform priors in log-space. As for the second visit, we cut the first 2.1 hours of the TSO and consider solely a linear slope for the systematics model. The transit light curves are then modelled in ExoTEP using the batman (Kreidberg, 2015) python package, and we explore the parameter space using and the Markov chain Monte Carlo (MCMC) sampler emcee (Foreman-Mackey et al., 2013). We run the MCMC for 10,000 steps and discard the first 6,000 steps as burn-in. The same procedure is repeated for the spectroscopic light curves (assuming the same systematics model and GP for each visit), where we have fixed the orbital parameters of TRAPPIST-1 c to the best-fit values from the white-light curve of the first visit ( = 28.55, = 0.112). The light curves are fit using 100 and 30 equal-pixel bins for the first and second spectral orders, respectively. The resulting spectra, binned to a fixed resolving power of , are shown in Figure 6.
Appendix B Additional Tables
@@
| Parameter | Value | ||
| Visit 1 | Visit 2 | Weighted Average | |
| T0 [BJD] | 2459881.401521 | 2460249.515096 | - |
| 0.133 | 0.136 | ||
| 28.31 | 28.12 | ||
Notes: For each visit, and the limb darkening parameters ( and ) were fit separately for each order. All other parameters were jointly fit to both.
@@
| 0.72 | 0.12 | 6050.99 473.23 | 6982.25 418.63 |
| 0.90 | 0.04 | 6585.89 111.83 | 6757.96 135.93 |
| 0.98 | 0.04 | 6772.63 91.97 | 6808.19 128.63 |
| 1.05 | 0.04 | 6881.03 71.42 | 6960.16 89.30 |
| 1.12 | 0.04 | 6820.05 69.73 | 6849.88 64.50 |
| 1.20 | 0.04 | 6699.58 63.55 | 6907.84 80.39 |
| 1.28 | 0.04 | 6694.65 64.58 | 6697.34 90.68 |
| 1.35 | 0.04 | 6971.35 60.31 | 7023.65 58.31 |
| 1.43 | 0.04 | 6886.42 76.10 | 7001.35 76.95 |
| 1.51 | 0.04 | 6847.72 84.78 | 6937.97 88.64 |
| 1.59 | 0.04 | 6789.90 76.61 | 6902.63 90.38 |
| 1.66 | 0.04 | 6895.75 74.07 | 7088.25 92.70 |
| 1.74 | 0.04 | 6914.99 90.90 | 6899.64 100.01 |
| 1.82 | 0.04 | 7006.82 97.16 | 7136.91 144.17 |
| 1.90 | 0.04 | 6804.84 126.15 | 6967.40 122.15 |
| 1.98 | 0.04 | 7032.96 118.65 | 7127.79 126.21 |
| 2.06 | 0.04 | 6796.19 132.58 | 6970.70 136.08 |
| 2.14 | 0.04 | 6842.86 136.57 | 6972.46 134.27 |
| 2.21 | 0.04 | 7010.63 135.26 | 6977.26 147.22 |
| 2.29 | 0.04 | 6919.63 128.65 | 6890.04 155.09 |
| 2.37 | 0.04 | 6707.69 153.51 | 6972.92 152.57 |
| 2.45 | 0.04 | 6643.73 180.88 | 6821.97 202.91 |
| 2.53 | 0.04 | 6770.44 232.33 | 7296.84 270.63 |
| 2.61 | 0.04 | 6837.59 233.79 | 6966.78 293.63 |
| 2.69 | 0.04 | 6183.53 235.11 | 6961.30 273.61 |
| 2.77 | 0.04 | 6589.56 270.29 | 6630.95 311.46 |

### v3

@@ 5.4 Implications for the Presence of an Atmosphere on TRAPPIST-1c
Although it could still be explained by some high metallicity scenarios, our non-detection of any significant atmospheric signature in the transmission spectrum of TRAPPIST-1 c is also in line with modeling predictions that TRAPPIST-1 c could be a bare rock, as any substantial atmosphere is expected to be stripped over the 5–9 Gyr lifetime of its host star (Burgasser & Mamajek, 2017), regardless of its initial volatile content (Krissansen-Totton, 2023).
@@
Even if they remain tentative (at the 2 level), our atmospheric retrieval results disfavor thick 1 bar atmospheres, even for high MMW compositions. The thin atmosphere scenarios that remain consistent with our observations given their smaller spectral features are however unlikely to be retained on TRAPPIST-1 c from a theoretical perspective. Even if Jeans escape (generally less efficient than blow-off) is expected to dominate for high-MMW atmospheres atop TRAPPIST-1 c, the escape rates even for resilient N2-CO2 atmospheres range from one Earth atmosphere per Myr to one Earth atmosphere mass per Gyr (Van Looveren et al., 2024) at the present-day irradiation of this planet. The implications are even stricter when considering the fact that any atmosphere on TRAPPIST-1 c also would have had to withstand the early XUV-active stages of the late M dwarf’s evolution (Bolmont et al., 2017), and its position inside of the runaway greenhouse limit (Turbet et al., 2020) indicates that water could not have been efficiently shielded from escape via condensation in a liquid water ocean, even if this limit is pushed to closer host star-planet distances around M dwarfs in 3D models (Turbet et al., 2023).
5.5 Impacts of Stellar Contamination on the Transmission Spectra
Our two-transit study of TRAPPIST-1 c aptly summarizes the promise and peril of transmission observations of rocky planets around M dwarf host stars. There are now a plethora of studies that have shown that stellar contamination is a true bottleneck in the attempt to detect atmospheres around rocky, M dwarf planets via transmission spectroscopy (e.g., Zhang et al., 2018; Wakeford et al., 2019; May et al., 2023; Moran et al., 2023; Lim et al., 2023; Cadieux et al., 2024; Rackham & de Wit, 2024). In most cases, different manifestations of the TLSE (i.e., different temperature distributions, covering fractions, etc. of stellar spots and faculae) give rise to obviously different transmission spectra between visits, allowing for a clear inference of the impacts of the TLSE, even without undertaking in-depth modelling. Consistency in the location and shapes of spectral features between visits has been argued to lend credence to an interpretation favouring a planetary atmosphere origin (e.g. May et al., 2023).
However, despite the high level of consistency between the transmission spectra (1.1- on average), our retrieval analysis still finds that both visits can be entirely explained by stellar contamination. As shown in Figure 8, and summarized in Table 2, we find very similar stellar heterogeneity parameters (spot/faculae temperatures and covering fractions) for both visits. We also find both visits to be spot-dominated, with minimal contributions from faculae, whereas Lim et al. (2023) have one spot-dominated, and one facula-dominated visit in their two-transit analysis of TRAPPIST-1 b.
The consistency between visits would perhaps be unsurprising if they were separated by a very small period of time, which would therefore probe similar spot and facula distributions on the host star. However, our two visits are separated by 367 d, significantly longer than the 3.3 d stellar rotation period (Luger et al., 2017). When phase-folded to the stellar rotation period, this corresponds to a phase difference of 0.21 — similar to the observations of Lim et al. (2023), who, as mentioned above, found no such consistency between their two visits. Moreover, both of our visits are at comparable stellar phases to those of Lim et al. (2023). This highlights the fact that the distributions of spots and faculae in the host star’s photosphere are entirely stochastic, and therefore TLSE models are non-predictive of future stellar contamination.
Irrespective of whether we perform TLS-only or joint TLS-atmosphere retrievals, we still infer consistent stellar contamination parameters (Figure 8). This is consistent with our finding that the retrieval does not pick up on any “true” atmosphere signature that would warrant changes in the inferred TLS parameters.
Finally, we evaluated the limiting effect of having to marginalize over stellar contamination on our sensitivity to thin 100% H2O atmospheres. This choice was motivated by the fact that water is the molecule most affected by this marginalization, given that its features coincide with prominent unocculted stellar spot signatures, but refer the reader to Piaulet-Ghorayeb et al. (in prep) for a more detailed exploration of the impact of TLSE marginalization on other infrared absorbers. We performed a retrieval on the TLS-corrected spectra from both visits (see Section 3.1), using the exact same setup as the joint atmosphere + TLS retrieval except that only the atmosphere properties are retrieved.
We find that our sensitivity drops by about two orders of magnitude through accounting for stellar contamination. As shown in Figure 9, without stellar contamination we could rule out at the 2 (3) level 100% H2O atmospheres with surface pressures down to 1 mbar (1.6 bar). However, when marginalizing over the TLSE, we can only rule out at the 2 (3) level cases with surface pressures of less than 0.1 bar (45 bar). Therefore, we are in a regime where the unknown TLS contribution to the spectrum, rather than the precision of the measured transit depths themselves, is the limiting factor in our sensitivity to the presence of high-MMW atmospheres. Furthermore, this effect will be even more problematic in other datasets where one needs to marginalize over different stellar contamination realizations for each visit in joint retrievals (e.g., Piaulet-Ghorayeb et al., in prep).
@@ 6 Conclusions
Here, we presented the first JWST transmission spectrum of the second innermost planet in the TRAPPIST-1 system; TRAPPIST-1 c. Our 0.6 – 2.85 µm NIRISS/SOSS spectrum reveals strong contamination by the TLSE in both visits. However, unlike in previous multi-visit analyses of rocky planets around M dwarfs (e.g., May et al., 2023; Moran et al., 2023; Cadieux et al., 2024), including TRAPPIST-1 b (Lim et al., 2023), we find consistent stellar heterogeneity properties between our two visits. Via an analysis of the in-transit data (i.e., the TLSE) and direct modelling of the out-of-transit stellar spectrum, we find the stellar photosphere to be dominated by spots 300 K cooler than the ambient photosphere temperature and with a covering fraction of 25%. Faculae have negligible contributions to both visits, irrespective of the analysis technique or choice of stellar models. This results in nearly identical manifestations of the TLSE and thus transmission spectra. Our results support the need for joint modelling of the TLSE and atmosphere properties to achieve robust constraints on rocky, M dwarf planets in the era of JWST. Importantly, further stellar model fidelity is critically required to accurately model out the impact of the TLSE and reach the sensitivity to detect or rule out thin atmospheres on small planets via their transit spectra.
When accounting for stellar contamination in our atmosphere analysis, we find that thick, clear hydrogen-dominated compositions are ruled out at better than 3- significance. The level of precision we reach does not enable us to rule out the cases with the highest MMW, where the atmosphere would be composed of 100% of a heavy volatile species such as H2O, NH3, or CH4, but even these scenarios are disfavored for 1 bar atmospheres at the -level due to the flatness of the TLSE-corrected spectrum. The sensitivity increases to partial pressures of about 10 mbar at 2 for the same species in atmospheres with N2 or O2 as the background gas.
@@ Appendix A NAMELESS Reduction and ExoTEP Light Curve Fitting
We reduce our two JWST NIRISS/SOSS visits of TRAPPIST-1c, starting from the uncalibrated raw data up to the extraction of the stellar spectra, using the NAMELESS pipeline (Coulombe et al., 2023; Feinstein et al., 2023). We first go through the Stages 1 and 2 steps of the STScI jwst pipeline (v1.12.5) (Bushouse et al., 2022), which include super-bias subtraction, reference pixel correction, non-linearity correction, ramp-fitting, and flat-fielding. We then proceed with a sequence of custom steps to correct for remaining sources of noise such as bad pixels, non-uniform background, cosmic-rays, and 1/ following the methods described in Benneke et al. (2024). We flag bad pixels by looking for pixels whose second derivative are significant (7) outliers in the spatial domain. The counts of all flagged pixels are then corrected using the bicubic interpolation method of the scipy.interpolate.griddata function. We correct for the non-uniform background by individually scaling the two regions of the STScI model background444https://jwst-docs.stsci.edu/ that are separated by the sudden jump in background flux situated around spectral pixel x700. Any remaining cosmic rays are corrected by clipping any count that is more than four standard deviations away from the running median for all pixels. As for the 1/ noise correction, we follow the same method described in Coulombe et al. (2023) and Benneke et al. (2024) where we scale individually each column of the first and second spectral orders considering only pixels that are within a 30-pixel distance from the center of the traces to compute the 1/ noise. We then extract the stellar spectra from both spectral orders using a simple box aperture with a width of 36 pixels.
@@
We perform the broadband and spectroscopic fits of the light curves with the ExoTEP framework (Benneke et al., 2019a, b). For our fit of the white-light curve, which is obtained by summing all wavelengths of the first spectral order (0.85–2.85 m), we keep free the mid-transit time (, considering uniform priors spanning the time of the observations for each visit), planet-to-star radius ratio (, [0.01,0.2]), semi-major axis (, [15,45]), and the impact parameter (, [0,1]). We consider the quadratic limb-darkening law and fit directly for and assuming large uniform priors ([-3,3]) following Coulombe et al. (2024). For the first visit, we consider a quadratic trend for the systematics model. We also include in the fit a Matérn 3/2 Gaussian Process using the celerite python (Foreman-Mackey et al., 2017) to remove higher-frequency trends that are not corrected by our systematics model. The timescale and amplitude of the GP model are kept free assuming large uninformative uniform priors in log-space. As for the second visit, we cut the first 2.1 hours of the TSO and consider solely a linear slope for the systematics model. The transit light curves are then modelled in ExoTEP using the batman (Kreidberg, 2015) python package, and we explore the parameter space using and the Markov chain Monte Carlo (MCMC) sampler emcee (Foreman-Mackey et al., 2013). We run the MCMC for 10,000 steps and discard the first 6,000 steps as burn-in. The same procedure is repeated for the spectroscopic light curves (assuming the same systematics model and GP for each visit), where we have fixed the orbital parameters of TRAPPIST-1 c to the best-fit values from the white-light curve of the first visit ( = 28.55, = 0.112). The light curves are fit using 100 and 30 equal-pixel bins for the first and second spectral orders, respectively. The resulting spectra, binned to a fixed resolving power of , are shown in Figure 7.
Appendix B Additional Figure & Tables
@@
Parameter | Prior Range | Value | ||
Visit 1 | Visit 2 | Weighted Average | ||
T0 [BJD] | [T02 hr] | 2459881.401528 | 2460249.515031 | |
[0.01, 0.9] | 0.08346 | 0.08344 | ||
[0.01, 0.9] | 0.08364 | 0.08415 | ||
[70, 90] | 89.73 | 89.38 | ||
[5, 50] | 28.35 | 28.18 | ||
[-1, 1] | 0.405 | 0.283 | ||
[-1, 1] | 0.168 | 0.346 | ||
[-1, 1] | 0.671 | 0.302 | ||
[-1, 1] | 0.098 | 0.105 |
Note: For each visit, and the limb darkening parameters ( and ) were fit separately for each order. All other parameters were jointly fit to both.
Table 2 contains the best-fitting parameters from the three-component stellar models via both the in-transit (TLS) and out-of-transit analyses.
Parameter | Model | In-Transit (TLS) | Out-of-Transit | ||
Visit 1 | Visit 2 | Visit 1 | Visit 2 | ||
Teff [K] | PHOENIX | ||||
SPHINX | |||||
PHOENIX | |||||
SPHINX | |||||
[K] | PHOENIX | 2437 | 2432 | ||
SPHINX | |||||
PHOENIX | |||||
SPHINX | |||||
PHOENIX | |||||
SPHINX | |||||
[K] | PHOENIX | ||||
SPHINX | 3054 | 3133 | |||
PHOENIX | |||||
SPHINX | 0.07 | 0.05 | |||
PHOENIX | |||||
SPHINX |
Note: For the stctm fits, the values quoted for and are derived from the posterior samples of , , and , and the intervals reported for the heterogeneity are derived from the and samples.
@@
0.72 | 0.12 | 6495.76 463.23 | 7080.25 918.63 |
0.90 | 0.04 | 6940.96 162.46 | 6884.63 193.99 |
0.98 | 0.04 | 6863.36 112.06 | 7026.99 147.35 |
1.05 | 0.04 | 7008.37 98.91 | 7049.40 95.48 |
1.12 | 0.04 | 6999.78 83.35 | 6889.46 77.15 |
1.20 | 0.04 | 6952.18 71.99 | 6988.60 74.63 |
1.28 | 0.04 | 6957.36 70.95 | 6955.08 72.35 |
1.35 | 0.04 | 7205.70 81.75 | 7081.42 78.11 |
1.43 | 0.04 | 7093.85 98.13 | 7079.19 95.84 |
1.51 | 0.04 | 7103.09 94.84 | 7154.09 93.00 |
1.59 | 0.04 | 7066.62 97.78 | 6949.11 120.23 |
1.66 | 0.04 | 7216.39 96.84 | 7171.97 95.48 |
1.74 | 0.04 | 7195.63 109.50 | 7011.15 110.87 |
1.82 | 0.04 | 7206.13 127.63 | 7295.62 123.62 |
1.90 | 0.04 | 7057.07 145.74 | 7217.11 158.78 |
1.98 | 0.04 | 7342.25 148.19 | 6961.31 144.52 |
2.06 | 0.04 | 7004.14 175.93 | 7073.29 167.36 |
2.14 | 0.04 | 7160.88 151.13 | 7154.43 157.82 |
2.21 | 0.04 | 7394.33 166.09 | 6897.93 147.24 |
2.29 | 0.04 | 6970.20 153.40 | 6997.21 172.76 |
2.37 | 0.04 | 6874.09 223.18 | 6547.71 198.38 |
2.45 | 0.04 | 7077.64 225.76 | 7116.19 245.70 |
2.53 | 0.04 | 7046.87 340.02 | 7388.53 327.30 |
2.61 | 0.04 | 7468.39 389.34 | 7336.15 420.91 |
2.69 | 0.04 | 6707.51 385.19 | 6884.10 356.58 |
2.77 | 0.04 | 6621.37 414.79 | 6666.19 417.26 |
We compare the posterior distributions of the spot, facula, and photosphere properties inferred from the TLS-only fit to the visit 1 and visit 2 spectra, and the joint atmosphere and stellar contamination retrievals (for the single-component atmosphere cases) in Figure 8.
Finally, we also present the results of our sensitivity study assessing the impact of marginalizing over the TLS effect on the atmosphere thicknesses that can be ruled out by our spectrum, in Figure 9.